Basic Salary of an employee is Rs. 18,000 and Allowances are Rs. 9000. According to the company’s policy casual leaves are 12 days per year and earned leaves are 24 days per year and normal working days are 22 per month.
Find the total cost of leaves as percent of gross salary.
1
Expert's answer
2019-05-16T10:58:19-0400

Total salary for the year is: (18,000 + 9,000)×12 = Rs. 324, 000.

Total number of working days is:

22×12 = 264 days.

The total cost of leaves as percent of gross salary is: (24 + 12)/264×100% = 13.64%.